a compound containing only C, H, and O, was extracted from the bark of the sassafrass tree. The combustion of 27.9 mg produced 75.7 mg of CO2 and 15.5 mg of CO2. The Molar mas of the compound was 162g/mol. Determine its empirical and molecular formula.
Answer #1

75.7 mg of CO2 and 15.5 mg of H2O.

mass of C = 12 * 75.7 / 44 = 20.65 mg

and

mass of H = 2 * 15.5 / 18 = 1.72 mg

mass of O = 27.9 - (20.65 + 1.72) = 5.53 mg

mole ratio:

C : H : O

20.65 / 12 : 1.72 / 1.0 : 5.53 / 16

or

1.72 : 1.72 : 0.346

or

1.72 / 0.346 : 1.72 / 0.346 : 0.346 / 0.346

or

5 : 5 : 1

thus

empirical formula is C5H5O

let molecular formula is (C5H5O)n

thus

n (5 * 12 + 5 * 1.0 + 1 * 16) = 162

or

n = 2

thus

molecular formula is C10H10O2.